Calculate the boiling point of a solution containing 3.5 g sugar (molar mass = 342 g)

dissolved in 150 g of water. Kb for water is 0.52 oC/m.




1
Expert's answer
2021-11-22T10:41:01-0500

Moles of sugar "=\frac{3.5}{342}=0.01moles"


Molarity "=\frac{0.01}{0.150}=0.067M"


"\u2206T_b=0.067\u00d70.52=0.03\u00b0c"


Boiling point "=100+0.03=100.03\u00b0c"